OpenStudy (anonymous):

Find (d^2y)/dx^2 if x^2+y^2=25 at x=4.

OpenStudy (ash2326):

@lizlozada did you try this?

OpenStudy (anonymous):

try what?

OpenStudy (ash2326):

this question

OpenStudy (anonymous):

yes, but I'm stuck

OpenStudy (ash2326):

ok, Would you mind sharing your work here? I'll guide you then

OpenStudy (anonymous):

2x+2yy'=0 (derivative of 25=0)

OpenStudy (anonymous):

THEN

OpenStudy (anonymous):

y'=-x/y what's next?

OpenStudy (ash2326):

Now we need to take the second derivative \[2x+2yy'=0\] \[\frac{d}{dx}(2x+2yy')=0\] \[2+2[y'\times \frac {d}{dx} (y)+2y y'']=0\] would you simplify this?

OpenStudy (anonymous):

why don't you take the derivative from -x/y?

OpenStudy (ash2326):

We could do that also. you try this. Then I'll take this way also

OpenStudy (anonymous):

i don't know how to do that though...

OpenStudy (ash2326):

It's easy. \[2+2[y'\times \frac {d}{dx} (y)+2y y'']=0\] you just have to take derivative of y and plugin here \[2+2[y'\times \large{\underline{\frac {d}{dx} (y)}}+2y y'']=0\] then just simplify it

OpenStudy (anonymous):

because when I asked my teacher, he said to take the derivative of -x/y, and I was curious as to how that would be done as well, if you don't mind.

OpenStudy (anonymous):

isn't the derivative of y, y'? so would you plug in y'?

OpenStudy (ash2326):

OK, we'll see both the methods \[2+2(y')^2+2yy''=0\] Now solve for y'' from here

OpenStudy (ash2326):

@lizlozada ??

OpenStudy (anonymous):

sorry i was helping my little sister with her essay

OpenStudy (ash2326):

no problem. Take you time

OpenStudy (anonymous):

isn't it -x/y because you factor a y' and divide by 2y?

OpenStudy (ash2326):

but the first term is 2. It doesn't have a y'

OpenStudy (anonymous):

oh right i didn't notice that. take way two from both sides then factor out a y'?

OpenStudy (ash2326):

Solve for y'' from here \[2+2(y')^2+2yy''=0\]

OpenStudy (anonymous):

so you don't take away 2 from both sides?

OpenStudy (ash2326):

Take 2 away, take y'' on one side

OpenStudy (anonymous):

then you factor out a y' from the left?

OpenStudy (ash2326):

\[1+(y')^2+y' y"=0\] \[y' y''=-(1+y')^2 \] divide both sides by y'

OpenStudy (ash2326):

Sorry this is the correct expression \[y' y''=-(1+y'^2)\]

OpenStudy (anonymous):

why is it 1?

OpenStudy (ash2326):

We have \[2+2(y')^2+2y' y''=0\] divide both sides by 2 \[1+(y')^2+y' y"=0\]

OpenStudy (anonymous):

oh ok. what's the difference between y' & y"?

OpenStudy (ash2326):

\[y'=\frac {dy}{dx}\] \[y''=\frac {d^2y}{dx^2}\]

OpenStudy (ash2326):

@lizlozada I'll be back in 10 minutes. You try this until then

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok i am very confused. would you mind teaching me how to take the derivative of -x/y? i feel that's simpler,

OpenStudy (ash2326):

Cool. Let's work that way

OpenStudy (anonymous):

thank you!

OpenStudy (ash2326):

we have \[\frac {dy}{dx}=\frac{-x}{y}\] Let's take derivative both sides, we get \[\frac{d^2y}{dx^2}=\frac{d}{dx} (\frac {-x}{y})\] do you know division rule of differentiation ?

OpenStudy (anonymous):

loDhi-hiDlo/ lolo?

OpenStudy (ash2326):

yes, apply that here

OpenStudy (anonymous):

y(1)-x(y')/y^2

OpenStudy (ash2326):

good, so we have \[\frac{d^2y}{dx^2}=\frac{y-xy'}{y^2}\] Let's find value of this at x=4, from x^2+y^2=25 what's value of y when x=4?

OpenStudy (anonymous):

which equation do i plug it into?

OpenStudy (ash2326):

x^2+y^2=25 this is the one which relates x and y

OpenStudy (anonymous):

then why do we get d^2y/dx^2 if we just put 4 in the og equation?

OpenStudy (ash2326):

we need to find the value of \(\frac{d^2y}{dx^2}\) at x=4, not value of y

OpenStudy (anonymous):

16+y^2=25 y^2=9 y=3

OpenStudy (anonymous):

then?

OpenStudy (ash2326):

Now find value of y' at x=4 and y=3

OpenStudy (ash2326):

\[\frac {dy}{dx}=\frac{-x}{y}\]

OpenStudy (anonymous):

-4/3

OpenStudy (ash2326):

Good :) Now substitute these values here to find d^2y/dx^2 \[\frac{d^2y}{dx^2}=\frac{y-xy'}{y^2}\]

OpenStudy (anonymous):

3-(4)(-4/3)/9 (4/3)/9 4/27?

OpenStudy (ash2326):

Check your calculations again :)

OpenStudy (anonymous):

ok sorry.(3+16/3)/9 =1

OpenStudy (ash2326):

\[\large\frac{3+\frac{16}{3}}9\] \[\large\frac{\frac{9+16}{3}}9\] \[\large\frac{\frac{25}{3}}9\] \[\large\frac{25}{27}\]

OpenStudy (ash2326):

@lizlozada read the whole post and do let me know if you have a doubt

OpenStudy (anonymous):

yes sorry I was rushing my math. everything is crystal clear, thank you so much.

OpenStudy (ash2326):

Are you certain?

OpenStudy (anonymous):

yes, i was thinking 16+9 is 27. so it would be 27/3=9, 9/9=1. i fully understand, thank you.

OpenStudy (ash2326):

Welcome :D